OpenStudy (anonymous):

hey can you see if I got this right 2. –6 + 3x = –9 (1 point)–1 –6 –5 –3 I think it is -5

OpenStudy (ajprincess):

is ur question? 6+3x=-9?

OpenStudy (anonymous):

4. –3x + 6 = –9 (1 point)–3 5 3 1 5. Paul rented a car for $129.00 plus $0.25 per mile. The total bill at the end of his trip was $216.50. Use the equation 129 + 0.25x = 216.50 to find the number of miles he drove. (1 point)1,382 miles 350 miles 864 miles 607 miles I think it is -3 and 607

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

ur answer -5 is right if ur question is 6+3x=-9

OpenStudy (anonymous):

it is how about the other ones

OpenStudy (ajprincess):

what is ur answer for 4th ques?

OpenStudy (anonymous):

for number four it is 5 and for the last one it is 350

OpenStudy (ajprincess):

ya 4th is right too:)

OpenStudy (ajprincess):

5th is too right:) very good:)